OpenStudy (anonymous):

Find all solutions in the interval [0, 2π). sec2 x - 2 = tan2 x

OpenStudy (anonymous):

@ganeshie8 i keep on getting no solution but i want to be sure

ganeshie8 (ganeshie8):

sec^2x -1 = tan^2x + 1 sec^2x - 1 = sec^2x -1 = 0 BCrap

ganeshie8 (ganeshie8):

so you're right ! no solutions :)

OpenStudy (anonymous):

yay thanks!

ganeshie8 (ganeshie8):

good job !!